OpenStudy (anonymous):

y<= –x + 7 y < 3x + 11

OpenStudy (anonymous):

How do I solve?

OpenStudy (anonymous):

solve for a variable and put it into the other

OpenStudy (turingtest):

as a system? or independently?

OpenStudy (anonymous):

As a system

OpenStudy (turingtest):

what imasofty said...

OpenStudy (turingtest):

wait, is one <= and the other just < ? or is that a typo?

OpenStudy (anonymous):

one is less then or equal to and the second one is just less than

OpenStudy (turingtest):

hm.... that throws me so I'm not sure, sorry. I just did this: solve y<= –x + 7 for x and plug it into the other inequality, which gave me y<8 but I think that may wrong because it doesn't take into account the <= sorry... I won't tell you I know if I'm not sure.

OpenStudy (anonymous):

yea theres somthing funny you gatta do...its from pre calc i think

OpenStudy (anonymous):

Im in alg II not precalc!

OpenStudy (turingtest):

moneybird is on the case, he should know what to do...

OpenStudy (anonymous):

thank you moneybird

OpenStudy (anonymous):

thank you everyone who is trying to help me!

OpenStudy (akshay_budhkar):

what i would have done is used a graph to solve it and got the answer on a graph paper, would have much simpler

OpenStudy (anonymous):

how do i graph it? y=mx+b?

OpenStudy (akshay_budhkar):

you can graph it by finding the intercepts on the x and y axis

OpenStudy (akshay_budhkar):

turning test will help you graph it he knows that :D lol turning you should have known this man u r better than me :D i need to go tc Sweden :D

OpenStudy (anonymous):

lol yes you do@ Akshay Budhkar... but im in USA now

OpenStudy (turingtest):

No, I think you and I are pretty close @akshay anyway, yes you can graph the individual lines like y=mx+b then use the inequality properties to find the solution area. I don't want to go through that step-by-step, but I'll see if I can get wolfram's answer: http://www.wolframalpha.com/input/?i=%5By%3C%3D+%E2%80%93x+%2B+7+%26%26+y+%3C+3x+%2B+11%5D

OpenStudy (turingtest):

ok, here's what I got all I see to do is to find the intersection of the two lines: 3x + 11= –x + 7 which is (-1,8) the intercepts of y=-x+7 are x=7 y=7 (that doesn't seem to matter here though) the x-intercept for y=3x+11 is -11/3 now all I can think to myself is that the graph looks something like this:...